\begin{thebibliography}{99}

\bibitem{and-ego-lan-tes}
K. Andreiev, I. Egorova, T.L. Lange and G. Teschl, 
{\it Rarefaction waves of 
the Korteweg-de Vries equation via nonlinear steepest deecent}, 
J. Differential Equations {\bf 10} (2016), pp. 5371-5410.
%\bibitem{barenblatt}
%G. I. Barenblatt, 
%{\it On the motion of suspended particles in a turbulent flow 
%taking up a half-space or a plane open channel of finite depth},
%Prikl. Mat. Meh. {\bf19} (1955), pp. 61-88 (in Russian).
%
%\bibitem{carillo-toscani}
%J. A. Carrillo and G. Toscani, 
%{\it Asymptotic $L^1$-decay of solutions of the porous medium equation 
%to self-similarity}, 
%Indiana Univ. Math. J. {\bf 49} (2000), pp. 113-142.

\bibitem{bona-schonbek}
J.L. Bona and M.E. Schonbek, 
{\it Travelling-wave solutions to the Korteweg-deVries-Burgers equation}, 
Proc. Roy. Soc. Edinburgh {\bf 101A} (1985), pp. 207-226. 

\bibitem{bona-rajopadhye-schonbek}
J.L. Bona, S.V. Rajopadhye and M.E. Schonbek, 
{\it Models for the propagation of bores I. Two dimensional theory}, 
Differential Integral Equations {\bf 7} (1994), pp. 699-734.

\bibitem{chh1}
R.P. Chhabra, 
{\it Bubbles, drops and particles in non-Newtonian Fluids}, 
CRC, Boca Raton, FL (2006).

\bibitem{chh2}
R.P. Chhabra, 
{\it Non-Newtonian Fluids: An Introduction}, 
URL http://www.physics.iitm.ac.in/$\tilde{\: }$compflu/Lect-notes/chhabra.pdf.

\bibitem{chh-ric}
R.P. Chhabra and J.F. Richardson,  
{\it Non-Newtonian flow and applied rheology}, 
2nd edn. Butterworth-Heinemann, Oxford (2008).

%\bibitem{de waele} 
%A. de Waele, 
%{\it Viscometry and plastometry}, 
%J. Oil Colour Chem. Assoc. {\bf 6} (1923), pp. 33-69.

\bibitem{du-gu} 
Q. Du and M.D. Gunzburger, 
{\it Analysis of a Ladyzhenskaya model for incompressible viscous flow}, 
J. Math. Anal. Appl. {\bf 155} (1991), pp. 21-45.

\bibitem{dua-fan-kim-xie}
R. Duan, L.-L. Fan, J.-S. Kim and L.-Q. Xie, 
{\it Nonlinear stability of strong rarefaction waves 
for the generalized KdV-Burgers-Kuramoto equation 
with large initial perturbation},
Nonlinear Anal. TMA {\bf 73} (2010), pp. 3254-3267.

\bibitem{dua-zha}
R. Duan and H.-J. Zhao, 
{\it Global stability of strong rarefaction waves 
for the generalized KdV-Burgers equation},
Nonlinear Anal. TMA {\bf 66} (2007), pp. 1100-1117.

\bibitem{ego-gru-tes}
I. Egorova, K. Grunert and G. Teschl, 
{\it On the Cauchy problem for 
the Korteweg-de Vries equation with steplike finite-gap initial data I. 
Schwartz-type perturbations}, 
Nonlinearity {\bf 22} (2009), pp. 1431-1457. 

\bibitem{ego-tes}
I. Egorova and G. Teschl, 
{\it On the Cauchy problem for 
the Korteweg-de Vries equation with steplike finite-gap initial data I\hspace{-.1em}I. 
Perturbations with finite moments}, 
J. Anal. Math. {\bf 115} (2011), pp. 71-101. 

\bibitem{gur-mac} 
M.E. Gurtin and R.C. MacCamy, 
{\it On the diffusion of biological populations}, 
Math. Biosci. {\bf 33} (1979), pp. 35-49.

\bibitem{harabetian}
E. Harabetian, 
{\it Rarefaction and large time behavior for parabolic 
equations and monotone schemes}, 
Comm. Math. Phys. {\bf 114} (1988), pp. 527-536.

\bibitem{hashimoto-matsumura} 
I. Hashimoto and A. Matsumura, 
{\it Large time behavior 
of solutions to an initial boundary value problem 
on the half space for scalar viscous conservation law}, 
Methods Appl. Anal. {\bf 14} (2007), pp. 45-59.

\bibitem{hattori-nishihara} 
Y. Hattori and K. Nishihara, 
{\it A note on the stability of rarefaction wave of the 
Burgers equation}, 
Japan J. Indust. Appl. Math. {\bf 8} (1991), pp. 85-96.
%
%\bibitem{huang-pan-wang} 
%F. Huang, R. Pan and Z. Wang, 
%{\it $L^1$Convergence to the Barenblatt solution for 
%compressible Euler equations with damping}, 
%Arch. Rational Mech. Anal. {\bf 200} (2011), pp. 665-689.

\bibitem{ilin-kalashnikov-oleinik} 
A.M. Il'in, A.S. Kala{\v{s}}nikov and O.A. Ole{\u\i}nik,
{\it Second-order linear equations of parabolic type}, 
Uspekhi Math. Nauk SSSR {\bf 17} (1962), pp. 3-146 (in Russian);
English translation in Russian Math. Surveys {\bf 17} (1962), pp. 1-143. 

\bibitem{ilin-oleinik} 
A.M. Il'in and O.A. Ole{\u\i}nik, 
{\it Asymptotic behavior of the solutions of the Cauchy problem for 
some quasi-linear equations for large values of the time}, 
Mat. Sb. {\bf 51} (1960), pp. 191-216 (in Russian).


\bibitem{jah-str-mul} 
P. Jahangiri, R. Streblow and D. M\"{u}ller, 
{\it Simulation of Non-Newtonian Fluids using Modelica}, 
Proceedings of the 9th International Modelica Conference September 3-5, 
Munich, Germany, (2012), pp. 57-62. 
%
%\bibitem{kamin} 
%S. Kamin, 
%{\it Source-type solutions for equations of nonstationary filtration}, 
%J. Math. Anal. Appl. {\bf 64} (1978), pp. 263-276. 

%\bibitem{kanel} 
%Y. Kanel', 
%{\it On a model system of one-dimensional gas motion}, 
%Differencial'nya Uravrenija {\bf 4} (1968), pp. 374-380. 

%\bibitem{kato1} 
%T. Kato, 
%{\it Linear evolution equations of ``hyperbolic tyle'' I}, 
%J. Fac. Sci. Univ. Tokyo Sect. A, {\bf 17} (1970), pp. 241-258. 
%
%\bibitem{kato2} 
%T. Kato, 
%{\it Linear evolution equations of ``hyperbolic tyle'' I, I\hspace{-.1em}I}, 
%J. Math. Soc. Japan, {\bf 19} (1973), pp. 648-666.
%
%\bibitem{kaw-matsu} 
%S. Kawashima and A. Matsumura, 
%{\it Stability of shock profiles in viscoelasticity 
%with non-convex constitutive relations}, 
%Comm. Pure Appl. Math. {\bf 47} (1994), pp. 1547-1569.

%\bibitem{lad1} 
%O.A. Lady{\v{z}}enskaja, 
%{\it New Equations for the Description of the Viscous Incompressible
%Fluids and Solvability in the Large of the Boundary Value Problems for Them, in 
%``Boundary Value Problems of Mathematical Physics V''}, 
%Amer. Math. Soc., 
%Providence, Rhode. Island, 1970.

%\bibitem{lad-sol-ura} 
%O. A. Lady{\v{z}}enskaja, V. A. Solonnikov and N. N. Ural'ceva, 
%{\it Linear and quasilinear equations of parabolic type}, 
%Transl. Math. Monographs, vol. 23, Amer. Math. Soc., 
%Providence, Rhode. Island, 1968.

\bibitem{lax} 
P.D. Lax, 
{\it Hyperbolic systems of conservation laws  I\hspace{-.1em}I}, 
Comm. Pure Appl. Math. {\bf 10} (1957), pp. 537-566.

\bibitem{liep-rosh} 
H.W. Liepmann and A. Roshko, 
{\it Elements of Gas Dynamics}, 
John Wiley \& Sons, Inc., New York, 1957.
%
%\bibitem{lions} 
%J. L. Lions, 
%{\it Quelques m{\`{e}}thodes de r{\`{e}}solution 
%des probl{\`{e}}mes aux limites non lin{\`{e}}aires}, 
%Dunod Gauthier-Villars, Paris, 1969 (in French).

\bibitem{liu-matsumura-nishihara} 
T.-P. Liu, A. Matsumura and K. Nishihara, 
{\it Behaviors of solutions for the Burgers equation 
with boundary corresponding to rarefaction waves}, 
SIAM J. Math. Anal. {\bf 29} (1998), pp. 293-308.

\bibitem{ma} 
J. M{\'{a}}lek, 
{\it Some frequently used models for non-Newtonian fluids}, 
URL http://www.karlin.mff.cuni.cz/$\tilde{\: }$malek/new/images/Lecture4.pdf.

\bibitem{ma-pr-st} 
J. M{\'{a}}lek, D. Pra{\v{z}}{\'{a}}k and M. Steinhauer,
{\it On the existence and regularity of solutions 
for degenerate power-law fluids}, 
Differential Integral Equations {\bf 19} (2006), pp. 449-462.
%
%\bibitem{matsu-mei} 
%A. Matsumura and M. Mei,
%{\it Nonlinear stability of viscous shock profile 
%for a non-convex system of viscoelasticity}, 
%Osaka J. Math. {\bf 34} (1997), pp. 589-603.

\bibitem{matsu-nishi1} 
A. Matsumura and K. Nishihara,
{\it Asymptotic toward the rarefaction wave of solutions of a
one-dimensional model system for compressible viscous gas}, Japan
J. Appl. Math. {\bf 3} (1986), pp. 1-13.

\bibitem{matsu-nishi2} 
A. Matsumura and K. Nishihara, 
{\it Asymptotics toward the rarefaction wave 
of the solutions of Burgers' equation with 
nonlinear degenerate viscosity
}
Nonlinear Anal. TMA {\bf 23} (1994), pp. 605-614.

\bibitem{matsu-nishi3} 
A. Matsumura and K. Nishihara, 
{\it Asymptotic stability of traveling waves for scalar 
viscous conservation laws with non-convex nonlinearity}, 
Comm. Math. Phys. {\bf 165} (1994), pp. 83-96.

\bibitem{matsumura-yoshida} 
A. Matsumura and N. Yoshida, 
{\it Asymptotic behavior of solutions to the Cauchy problem 
for the scalar viscous conservation law 
with partially linearly degenerate flux}, 
SIAM J. Math. Anal. {\bf 44} (2012), pp. 2526-2544. 

\bibitem{matsumura-yoshida'} 
A. Matsumura and N. Yoshida, 
{\it Global asymptotics toward 
the rarefaction waves for solutions to the Cauchy problem of
the scalar conservation law with nonlinear viscosity}, 
Osaka J. Math. {\bf 57} (2020), pp. 187-205.

\bibitem{nishi-raj}
K. Nishihara and S.V. Rajopadhye, 
{\it Asymptotic behavior of solutions to the Korteweg-deVries-Burgers equation}, 
Differential Integral Equations {\bf 11} (1998), pp. 85-93.
%
%\bibitem{nagai-mimura1} 
%T. Nagai and M. Mimura, 
%{\it Some nonlinear degenerate diffusion equations 
%related to population dynamics}, 
%J. Math. Soc. Japan {\bf 35} (1983), pp. 539-562.
%
%\bibitem{nagai-mimura2} 
%T. Nagai and M. Mimura, 
%{\it Asymptotic behavior for 
%a nonlinear degenerate diffusion equation in population dynamics}, 
%SIAM J. Appl. Math. {\bf 43} (1983), pp. 449-464.

%\bibitem{ost} 
%W. Ostwald, 
%{\it \"{U}ber die Geschwindigkeitsfunktion der Viskositat disperser Systeme}, 
%I. Colloid Polym. Sci. {\bf 36} (1925), pp. 99-117 (in German).

%
%\bibitem{pattle} 
%R. E. Pattle, 
%{\it Diffusion from an instantaneous point source with 
%a concentration-dependent coefficient}, 
%Quart. J. Mech. Appl. Math. {\bf 12} (1959), pp. 407-409. 

\bibitem{rajopadhye}
S.V. Rajopadhye, 
{\it Decay rates for the solutions of model equations for bore propagation}, 
Proc. Roy. Soc. Edinburgh {\bf 125A} (1995), pp. 371-398. 

\bibitem{rua-gao-che}
L.-Z. Ruan, W.-L. Gao and J. Chen, 
{\it Asymptotic stability of the rarefaction wave 
for the generalized KdV-Burgers-Kuramoto equation},
Nonlinear Anal. TMA {\bf 68} (2008), pp. 402-411.

\bibitem{smoller} 
J. Smoller, 
{\it Shock Waves and Reaction-diffusion Equations}, Springer-Verlag, 
New York-Berlin, 1983. 

\bibitem{soc} 
T. Sochi, 
{\it Pore-Scale Modeling of Non-Newtonian Flow in Porous Media}, 
PhD thesis, Imperial College London, 2007. 
%
%\bibitem{vaz1} 
%J. L. V{\'{a}}zquez, 
%{\it Smoothing and Decay Estimates for Nonlinear Diffusion Equations: 
%Equations of Porous Medium Type}, Oxford Math. and Appl., 2006. 
%
%\bibitem{vaz2} 
%J. L. V{\'{a}}zquez, 
%{\it The Porous Medium Equation: Mathematical Theory}, 
%Oxford Math. Monogr., 2007. 

\bibitem{wan-zhu} 
Z.-A. Wang and C.-J. Zhu, 
{\it Stability of the rarefaction wave 
for the generalized KdV-Burgers equation},
Acta Math. Sci. {\bf 22B}(3) (2002), pp. 319-328.

\bibitem{yoshida1} 
N. Yoshida, 
{\it Decay properties of solutions toward a multiwave pattern 
for the scalar viscous conservation law 
with partially linearly degenerate flux}, 
Nonlinear Anal. TMA {\bf 96} (2014), pp. 189-210.

\bibitem{yoshida2} 
N. Yoshida, 
{\it Decay properties of solutions 
to the Cauchy problem for the scalar conservation law 
with nonlinearly degenerate viscosity}, 
Nonlinear Anal. TMA {\bf 128} (2015), pp. 48-76.

\bibitem{yoshida3} 
N. Yoshida, 
{\it Asymptotic behavior of solutions toward a multiwave pattern 
for the scalar conservation law 
with the Ostwald-de Waele-type viscosity}, 
SIAM J. Math. Anal. {\bf 49} (2017), pp. 2009-2036.

\bibitem{yoshida4} 
N. Yoshida, 
{\it Decay properties of solutions toward a multiwave pattern 
to the Cauchy problem for the scalar conservation law 
with degenerate flux and viscosity}, 
J. Differential Equations {\bf 263} (2017), pp. 7513-7558.

\bibitem{yoshida5} 
N. Yoshida, 
{\it Asymptotic behavior of solutions toward the viscous shock waves 
to the Cauchy problem 
for the scalar conservation law with nonlinear flux and viscosity}, 
SIAM J. Math. Anal. {\bf 50} (2018), pp. 891-932.

\bibitem{yoshida6} 
N. Yoshida, 
{\it Asymptotic behavior of solutions toward 
the rarefaction waves to the Cauchy problem for 
the scalar diffusive dispersive conservation laws}, 
Nonlinear Anal. {\bf 189} (2019), pp. 1-19.

\bibitem{yoshida7} 
N. Yoshida, 
{\it Global structure of solutions toward 
the rarefaction waves for the Cauchy problem 
of the scalar conservation law with nonlinear viscosity}, 
J. Differential Equations {\bf 269} (2020), pp. 10350-10394.

\bibitem{yoshida8} 
N. Yoshida, 
{\it Asymptotic behavior of solutions toward a multiwave pattern 
to the Cauchy problem for the dissipative wave equation 
with partially linearly degenerate flux}, 
Funkcialaj Ekvacioj {\bf 64} (2021), pp. 49-73.

%
%\bibitem{zel-kom} 
%Ya. B. Zel'dovi{\v{c}} and A. S. Kompaneec, 
%{\it On the theory of propagation 
%of heat with the heat conductivity depending upon the temperature}, 
%Collection in honor of the seventieth birthday of academician A. F. Ioffe, 
%Izdat. Akad. Nauk SSSR, (1950), pp. 61-71 (in Russian). 

\end{thebibliography}

\documentclass[preprint,10pt]{elsarticle}

\usepackage{mathrsfs,amsmath,amssymb}

\usepackage{mathrsfs,amsmath,amssymb}
\usepackage{mathrsfs,amsmath,amssymb}
\usepackage{amscd} 
\usepackage{relsize,amsmath} 

\makeatletter
\newcommand{\biggg}[1]{{\hbox{$\left#1\vbox to 20.5pt{}\right.\n@space$}}}
\newcommand{\Biggg}[1]{{\hbox{$\left#1\vbox to 23.5pt{}\right.\n@space$}}}
\newcommand{\bigggg}[1]{{\hbox{$\left#1\vbox to 26.5pt{}\right.\n@space$}}}
\newcommand{\Bigggg}[1]{{\hbox{$\left#1\vbox to 29.5pt{}\right.\n@space$}}}
\newcommand{\biggggg}[1]{{\hbox{$\left#1\vbox to 32.5pt{}\right.\n@space$}}}
\newcommand{\Biggggg}[1]{{\hbox{$\left#1\vbox to 35.5pt{}\right.\n@space$}}}
\newcommand{\bigggggg}[1]{{\hbox{$\left#1\vbox to 38.5pt{}\right.\n@space$}}}
\newcommand{\Bigggggg}[1]{{\hbox{$\left#1\vbox to 41.5pt{}\right.\n@space$}}}
\makeatother

\makeatletter
\@addtoreset{equation}{section}
\makeatother

\usepackage{braket}

\begin{document}

\newtheorem{thm}{Theorem}
\newtheorem{lem}[thm]{Lemma}
\newdefinition{rmk}{Remark}
\newproof{pf}{Proof}
\newproof{pot}{Proof of Theorem \ref{thm2}}

\begin{frontmatter}



\title{Asymptotic behavior of solutions toward the constant state 
to the Cauchy problem for the non-viscous diffusive dispersive conservation law}


\author[labe1
]{Natsumi Yoshida}
\ead{14v00067@gmail.com
}

\address[label]{Graduate Faculty of Interdisciplinary Research Faculty of Education, 
University of Yamanashi, Kofu, Yamanashi 400-8510, Japan.}
\address{}

\begin{abstract}
In this paper, we investigate the asymptotic behavior of solutions 
to the Cauchy problem 
for the scalar non-viscous diffusive dispersive conservation laws 
where the far field states are prescribed. 
We proved that 
the solution of the Cauchy problem tends toward the constant state 
as time goes to infinity.
\end{abstract}

\begin{keyword} 
diffusive dispersive conservation laws \sep non-viscous diffusive flux \sep convex flux 
\sep asymptotic behavior \sep constant state

\medskip
AMS subject classifications: 35K55, 35B40, 35L65
\end{keyword}

\end{frontmatter}

%



\pagestyle{myheadings}
\thispagestyle{plain}
\markboth{N. YOSHIDA}{NON-VISCOUS DIFFUSIVE DISPERSIVE CONSERVATION LAWS}

\section{Introduction and main theorems}
In this paper, 
we consider the asymptotic behavior of solutions to the Cauchy problem 
for a one-dimensional scalar diffusive dispersive conservation laws 
without viscous flux 
\begin{eqnarray}
 \left\{\begin{array}{ll}
  \partial_tu +\partial_x \bigl( \, 
f(u) + \delta \, \partial_x^2u 
+ \nu \, \partial_x^3u 
 \, \bigr) =0
  \qquad &\big( \, t>0, \: x\in \mathbb{R} \, \big), \\[5pt]
 u(0, x) = u_0(x) \rightarrow \tilde{u} \qquad &( x \rightarrow \pm \infty ),
 \end{array}
 \right.\,
\end{eqnarray}
where, $u=u(t, x)$ is the unknown function of $t>0$ and $x\in \mathbb{R}$, 
the so-called conserved quantity, 
$\tilde{u}\in \mathbb{R}$ is the constant state, 
$$
f(u) + \delta \, \partial_x^2u 
+ \nu \, \partial_x^3u 
\quad \big( \, \delta \in \mathbb{R}, \; \nu \geq 0 \, \big)
$$ 
is the total flux 
(that is, the functions $f(u)$, 
$\delta \, \partial_x^2u$ 
and $\nu \, \partial_x^3u$ stand for 
the convective flux, 
dispersive one 
and diffusive one, respectively), 
$u_0$ is the initial data, 
and $u_{\pm } \in \mathbb{R}$ 
are the prescribed far field states. 
We suppose that $f$ is a smooth function. 
It is noted that, the equation in the problem (1.1) is the non-viscous case ($\mu=0$) for the 
following equation:
\begin{equation}
\partial_tu +\partial_x \bigl( \, 
f(u) - \mu \, \partial_xu + \delta \, \partial_x^2u 
+ \nu \, \partial_x^3u 
 \, \bigr) =0, 
\end{equation} 
where $\mu \, \partial_xu$ is viscous/diffusive flux 
($\mu$ is the so-called viscous coefficient or anti-diffusion coefficient). 
It should be noted that, in the case $\mu > 0$, $\delta =0$, $\nu = 0$, 
(1.2) becomes the viscous conservation law/generalized viscous Burgers equation, 
in the case $\mu = 0$, $\delta \in \mathbb{R}$, $\nu = 0$, 
the one does the Korteweg-de Vries equation as one of the dispersive conservation laws, 
in the case $\mu > 0$, $\delta \in \mathbb{R}$, $\nu = 0$, 
the one does the generalized Korteweg-de Vries-Burgers equation,
in the case $\mu > 0$, $\delta \in \mathbb{R}$, $\nu > 0 $, 
the one does the generalized Korteweg-de Vries-Burgers-Kuramoto equation or 
the derivative form of the Kuramoto-Sivashinsky equation. 
We also note that the Korteweg-de Vries equation
can be categorized as dispersive conservation laws,
and the Korteweg-de Vries-Burgers equation 
and the Korteweg-de Vries-Burgers-Kuramoto equation or 
the derivative form of the Kuramoto-Sivashinsky equation the 
diffusive dispersive conservation laws. 

There have been known the various of the stability results 
concerning with the conservation laws (see \cite{and-ego-lan-tes}, \cite{bona-rajopadhye-schonbek}, 
\cite{du-gu}, \cite{dua-zha}, \cite{ego-gru-tes}, \cite{ego-tes}, 
\cite{harabetian}, \cite{hashimoto-matsumura}, \cite{hattori-nishihara}, \cite{ilin-oleinik}, 
\cite{lax}, \cite{liu-matsumura-nishihara}, 
\cite{matsu-nishi1}, \cite{matsu-nishi2}, \cite{matsu-nishi3}, 
\cite{matsumura-yoshida}, \cite{matsumura-yoshida'}, 
\cite{nishi-raj}, 
\cite{rajopadhye}, 
\cite{wan-zhu}, 
\cite{yoshida1}, \cite{yoshida2}, \cite{yoshida3}, \cite{yoshida4}, \cite{yoshida5}, 
\cite{yoshida7}, \cite{yoshida8} 
and so on, cf. \cite{chh1}, \cite{chh2}, 
\cite{chh-ric}, \cite{gur-mac}, \cite{ilin-kalashnikov-oleinik}, \cite{jah-str-mul}, 
\cite{liep-rosh}, 
\cite{ma}, \cite{ma-pr-st}, \cite{smoller}, \cite{soc}). 
In partucular, for the Cauchy problem of (1.2), 
Duan-Fan-Kim-Xie \cite{dua-fan-kim-xie}, 
Ruan-Gao-Chen \cite{rua-gao-che}, 
showed some stabilities of the rarefaction waves and 
Yoshida \cite{yoshida6} showed the global stabilities of the constant state 
and the rarefaction wave. 

However, the any stabilities have not yet been known for more difficult non-viscous case, (1.1). 

Our main theorem is stated as follows.  

\medskip

\noindent
{\bf Theorem 1.1} (Main Theorem ){\bf .}\quad{\it
Assume 
the convective flux $f \in C^2(\mathbb{R})$ satisfy 
\begin{equation}
|\, f''(u) \,| \leq O(1)\, \big( \, 1 + |\, u \,|^q \, \big)
\quad \big( \, 0\leq q\leq5 \, \big) 
\end{equation}
and the initial data satisfy
$u_0-\tilde{u} \in L^2$ and
$\partial _xu_0 \in H^1$. 
Then the Cauchy problem {\rm(1.1)} has a 
unique global in time 
solution $u$ 
satisfying 
\begin{eqnarray*}
 \left\{\begin{array}{ll}
u-\tilde{u} \in C^0\bigl( \, [\, 0, \, \infty \, ) \, ; H^2 \, \bigr), \\[5pt]
\partial _x^2 u \in L^2\bigl( \, 0,\, \infty \, ; H^2 \, \bigr),
\end{array} 
\right.\,
\end{eqnarray*}
and the asymptotic behavior 
$$
\lim _{t \to \infty}
\bigg( \,
\sup_{x\in \mathbb{R}}
| \, u(t,x) - \tilde{u} \, | 
+\sup_{x\in \mathbb{R}}
 |\,\partial_xu (t, x)\,| 
\, \bigg)= 0.
$$
}

\medskip

The proofs of Theorem 1.1 
is given 
by a technical energy method. 

\medskip

This paper is organized as follows. 
In Section 2, We reformulate the problem 
in terms of the deviation from 
the asymptotic state. 
In order to show the asymptotics, 
we establish the {\it a priori} estimates 
by using the technical energy method in Section 3.  

\medskip

{\bf Some Notation.}\quad 
We denote by $C$ generic positive constants unless 
they need to be distinguished. 
In particular, use 
$C_{\alpha, \beta, \cdots }$ 
when we emphasize the dependency on $\alpha,\: \beta,\: \cdots $.

For function spaces, 
${L}^p = {L}^p(\mathbb{R})$ and ${H}^k = {H}^k(\mathbb{R})$ 
denote the usual Lebesgue space and 
$k$-th order Sobolev space on the whole space $\mathbb{R}$ 
with norms $||\cdot||_{{L}^p}$ and $||\cdot||_{{H}^k}$, 
respectively. 

\bigskip 

\noindent
\section{Reformulation of the problem} 
In this section, we reformulate our problem (1.1) 
in terms of the deviation from the asymptotic state. 
Now letting 
\begin{equation}
u(t, x) = \tilde{u} + \psi(t, x), 
\end{equation}
we reformulate the problem (1.1) in terms of 
the deviation $\psi $ from $\tilde{u}$ as 
\begin{eqnarray}
 \left\{\begin{array}{ll}
  \partial _t\psi 
  + \partial_x \big( \, f(\psi+\tilde{u}) \, \big) 
  =-\delta \, \partial_x^3 \psi - \nu \, \partial_x^4 \psi 
     \quad  \bigl( \, t>0,\: x\in \mathbb{R} \, \bigr), \\[5pt]
  \psi(0, x) = \psi_0(x) 
  := u_0(x)-\tilde{u} %
  \rightarrow 0 \quad (x\rightarrow \pm \infty).
 \end{array}
 \right.\,
\end{eqnarray}
Then we look for 
the unique global in time solution 
$\psi $ which has the asymptotic behavior 
\begin{equation}
\displaystyle{
\sup_{x \in \mathbb{R}}|\,  \psi(t, x) \, |
\to  0, \quad 
\sup_{x\in \mathbb{R}}
 |\,\partial_x\psi (t, x)\,|\to  0 
\qquad (t\to  \infty)}. 
\end{equation}
Here we note 
that $\psi_0 \in H^2$.
Then the corresponding theorems 
for $\psi$ to Theorems 1.1 we should prove is stated as follows. 

\medskip

\noindent
{\bf Theorem 2.1.} (Global Existence){\bf .}\quad{\it
Assume 
the convective flux $f \in C^2(\mathbb{R})$ 
satisfy {\rm(1.3)} and the initial data satisfy
$\psi_0 \in H^2$. 
Then the Cauchy problem {\rm(2.2)} has a 
unique global in time 
solution $\psi$ 
satisfying 
\begin{eqnarray*}
\left\{\begin{array}{ll}
\psi \in C^0\bigl( \, [\, 0, \, \infty ) \, ; H^2 \, \bigr), \\[5pt]
\partial _x^2 \psi \in L^2\bigl( \, 0,\, \infty \, ; H^2 \, \bigr),
\end{array} 
\right.\,
\end{eqnarray*}
and the asymptotic behavior 
$$
\lim _{t \to \infty}
\bigg( \,
\sup_{x\in \mathbb{R}}
|\,\psi (t, x)\,| 
+\sup_{x\in \mathbb{R}}
 |\,\partial_x\psi (t, x)\,| 
\, \bigg)= 0. 
$$
}

%
%

\medskip

To accomplish the proofs of Theorem 2.1, 
we prepare the local existence precisely, 
we formulate 
the problem (2.2) at general 
initial time $\tau \ge 0$: 
\begin{eqnarray}
 \left\{\begin{array}{ll}
  \partial _t\psi 
  + \partial_x \left( \, f(\psi+\tilde{u}) \, \right) 
  = - \delta \, \partial_x^3 \psi - \nu \, \partial_x^4 \psi 
     \quad  \bigl( \, t>\tau,\: x\in \mathbb{R} \, \bigr), \\[5pt]
  \psi(\tau, x) = \psi_\tau(x) 
  := u_\tau(x)-\tilde{u} %
  \rightarrow 0 \quad (x\rightarrow \pm \infty).
 \end{array}
 \right.\,
\end{eqnarray}

Then the local existence theorem is stated as follows 
(since the proof is standard, 
we state only here (cf. \cite{dua-fan-kim-xie}, \cite{dua-zha}, 
\cite{rua-gao-che}, \cite{wan-zhu})). 

\medskip

\noindent
{\bf Theorem 2.2} (Local Existence){\bf .}\quad{\it
For any $ M > 0 $, there exists a positive constant 
$t_0=t_0(M)$ not depending on $\tau$ 
such that if 
$\psi_{\tau} \in H^2$ and 
$$
\displaystyle{ 
\| \, \psi_{\tau} \, \|_{L^2} + \| \, \partial_x^2\psi_{\tau} \, \|_{L^2} \leq M}, 
$$
then the Cauchy problem {\rm (2.4)} has a unique solution $\psi$ 
on the time interval $[\, \tau, \, \tau+t_{0}(M)\, ]$ satisfying 
\begin{eqnarray*}
\left\{\begin{array}{ll}
\psi \in C^0
\bigl( \, [\, \tau, \, \tau+t_{0}\, ] \, ; H^2 \bigr),\\[5pt]
\partial _x^2 \psi  \in L^2\bigl( \, \tau, \, \tau+t_{0} \, ; H^2 \bigr),\\[5pt]
\displaystyle{
\sup_{t \in [\tau, \tau + t_{0}]} \, 
\left( \, 
\| \, \psi(t) \, \|_{L^2} + \| \, \partial_x^2\psi(t) \, \|_{L^2}
\, \right)
}
\leq 2 \, M. 
\end{array} 
\right.\,
\end{eqnarray*}
}

\medskip

Next, we state the {\it a priori} estimates as follows. 

\medskip

\noindent
{\bf Theorem 2.3} ({\it A Priori} Estimates){\bf .}\quad{\it 
Under the same assumptions as in Theorem 2.1, 
for any initial data 
$\psi_0 \in H^2$, 
there exists a positive constant $C_{\psi_0}$
such that 
if the Cauchy problem {\rm (2.1)}
has a solution 
$\phi$ 
on the time interval $[\, 0, \, T\, ]$ satisfying 
\begin{eqnarray*}
\left\{\begin{array}{ll}
\phi \in C^0
\bigl( \, [\, 0, \, T\, ] \, ; H^2 \bigr),\\[5pt]
\partial _x^2 \psi  \in L^2\bigl( \, 0, \, T \, ; H^2 \bigr),
\end{array} 
\right.\,
\end{eqnarray*}
for some positive constant $T$, 
then it holds that 
\begin{align}
\begin{aligned}
&\| \, \psi(t) \, \|_{H^2}^{2} 
  + \int^{t}_{0} 
\| \, \partial_{x}^2\psi(\tau) \, \|_{H^2}^{2} \, \mathrm{d}\tau
+\int^{t}_{0} 
\left( \, 
\sup_{x \in \mathbb{R}} |\, \psi(\tau,x) \, |
\, \right)^8
\, \mathrm{d}\tau \\
&
+ \int^{t}_{0} 
\left( \, 
\sup_{x \in \mathbb{R}} |\, \partial_x\psi(\tau,x) \, |
\, \right)^{\frac{8}{3}}
\, \mathrm{d}\tau 
\leq C_{\psi_0} 
\quad \big( \, t \in [\, 0, \, T\, ] \, \big).
\end{aligned}
\end{align}
}

\medskip

Combining the local existence Theorem 2.2 together with 
the each {\it a priori} estimates, Theorem 2.3 , we can obtain 
global existence Theorem 2.1. 
In fact, we can obtain the unique global in time solutions $\psi$ to (2.2) in 
Theorem 2.2 satisfying 
\begin{eqnarray*}
\left\{\begin{array}{ll}
\psi \in C^0\bigl( \, [\, 0, \, \infty ) \, ; H^2 \, \bigr), \\[5pt]
\partial _x^2 \psi \in L^2\bigl( \, 0,\, \infty \, ; H^2 \, \bigr),
\end{array} 
\right.\,
\end{eqnarray*}
and 
\begin{equation}
\sup_{t \ge 0}
  \| \, \psi(t) \, \|_{H^2}^{2} 
+ \int^{\infty}_{0} 
\| \, \partial_{x}^2\psi(t) \, \|_{H^2}^{2} \, \mathrm{d}t
< \infty 
\end{equation}
which yields 
\begin{equation}
\int _0^{\infty }\bigg|\,\frac{\mathrm{d}}{\mathrm{d}t} 
\|  \, \partial_x^2\psi(t)  \, \|_{L^2}^2 \,\bigg| \, \mathrm{d}t
< \infty.
\end{equation}
%
%
%
%
We immediately have from (2.6) and (2.7) that 
\begin{equation}
\|  \, \partial_x\psi(t)  \, \|_{L^2} \to 0\quad (t \to \infty).
\end{equation}
Further from (2.8) with $T\rightarrow \infty$ and 
\begin{equation}
\sup_{x \in \mathbb{R}}|\,\partial_x\psi(t,x)\,|
\leq \sqrt{2}\, \| \,\partial_x\psi(t)\, \|^{\frac{1}{2}}_{L^2} 
\| \,\partial_x^2\psi(t)\, \| ^{\frac{1}{2}}_{L^2} 
\quad \big( \, t \ge 0 \, \big), 
\end{equation} 
we obtain the asymptotic behavior (2.3). 

Thus Theorem 2.1 is proved.


\bigskip 

\noindent
\section{{\it A priori} estimates}
In this section, under the assumption 
\begin{equation}
|\, f''(u) \,| \leq O(1)\, \big( \, 1 + |\, u \,|^q \, \big) \quad \big( \, q\ge0 \, \big), 
\end{equation}
we show the following {\it a priori} estimate for $\psi$ in Theorem 2.3. 
To do that, we prepare the following basic estimate. 

\medskip

\noindent
{\bf Proposition 3.1.}\quad {\it
For $q \ge 0$, it follows that 
\begin{equation*}
\| \, \psi(t) \, \|_{L^2}^{2} +
2\, \nu \,
\int^{t}_{0} 
\| \, \partial_{x}^2\psi(\tau) \, \|_{L^2}^{2} \, \mathrm{d}\tau 
=\| \, \psi_0 \, \|_{L^2}^{2}
\quad \big( \, t \in [\, 0, \, T\, ] \, \big). 
\end{equation*}
}

\medskip

{\bf Proof of Proposition 3.1.}
Multiplying the equation in (2.2) by $\phi$ and integrating it with respect to $x$, 
we have, after integration by parts, 
\begin{equation}
\frac{1}{2} \, \frac{\mathrm{d}}{\mathrm{d}t} \, \| \, \psi(t) \, \|_{L^2}^2 
+ \nu \, \| \, \partial_{x}^2\psi(t) \, \|_{L^2}^{2} 
=0.
\end{equation}
Next, integrating (3.2) with respect to $t$, we immediately get the desired estimate.

Thus, we complete the proof of Proposition 3.1. 

\medskip

From Proposition 3.1, we have the next lemma. 

\medskip

\noindent
{\bf Lemma 3.2.}\quad {\it
There exists a positive constant $C_{\psi_0}$
such that 
\begin{equation*}
\int^{t}_{0} 
\left( \, 
\sup_{x \in \mathbb{R}} |\, \psi(\tau,x) \, |
\, \right)^8
\, \mathrm{d}\tau 
+ \int^{t}_{0} 
\left( \, 
\sup_{x \in \mathbb{R}} |\, \partial_x\psi(\tau,x) \, |
\, \right)^{\frac{8}{3}}
\, \mathrm{d}\tau 
\leq C_{\phi_0} 
\quad \big( \, t \in [\, 0, \, T\, ] \, \big). 
\end{equation*}
}

\medskip

{\bf Proof of Lemma 3.2.}
By using the Sobolev inequality and the integration by parts, 
we get 
\begin{align}
\begin{aligned}
\sup_{x \in \mathbb{R}}|\,\psi(t,x)\,|
&\le \sqrt{2}\, \| \,\psi(t)\, \|^{\frac{1}{2}}_{L^2} 
\| \,\partial_x\psi(t)\, \| ^{\frac{1}{2}}_{L^2}\\
&\le \sqrt{2}\, \| \,\psi(t)\, \|^{\frac{3}{4}}_{L^2} 
\| \,\partial_x^2\psi(t)\, \| ^{\frac{1}{4}}_{L^2}
\quad \big( \, t \in [\, 0, \, T\, ] \, \big),
\end{aligned}
\end{align}
\begin{align}
\begin{aligned}
\sup_{x \in \mathbb{R}}|\,\psi(t,x)\,|
&\le \sqrt{2}\, \| \,\partial_x\psi(t)\, \|^{\frac{1}{2}}_{L^2} 
\| \,\partial_x^2\psi(t)\, \| ^{\frac{1}{2}}_{L^2}\\
&\le \sqrt{2}\, \| \,\psi(t)\, \|^{\frac{1}{4}}_{L^2} 
\| \,\partial_x^2\psi(t)\, \| ^{\frac{3}{4}}_{L^2}
\quad \big( \, t \in [\, 0, \, T\, ] \, \big).
\end{aligned}
\end{align}
From (3.3) and (3.4), noting Proposition 3.1, we immediately have the desired estimate.

Thus, the proof is complete.

\medskip

Next, we state the {\it a priori} estimate for $\partial_{x}^2\psi$ as follows. 

\medskip

\noindent
{\bf Proposition 3.3.}\quad {\it
For $0\le q \le 5$, there exists a positive constant $C_{\psi_0}$
such that 
\begin{equation*}
\| \, \partial_{x}^2\psi(t) \, \|_{L^2}^{2} 
  + \int^{t}_{0} 
\| \, \partial_{x}^4\psi(\tau) \, \|_{L^2}^{2} \, \mathrm{d}\tau
\leq C_{\psi_0} 
\quad \big( \, t \in [\, 0, \, T\, ] \, \big).
\end{equation*}
}

Once Proposition 3.3 holds true, by using Proposition 3.1, 
we can estimate as follows. 
\begin{equation}
\| \, \partial_{x}\psi \, \|_{L^2}^{2} 
\le \| \, \psi \, \|_{L^2} \, \| \, \partial_{x}^2\psi \, \|_{L^2} 
\le C_{\phi_0},
\end{equation}
\begin{equation}
\int^{t}_{0} 
\| \, \partial_{x}^3\psi \, \|_{L^2}^{2} \, \mathrm{d}\tau
\le \left( \, \int^{t}_{0} 
    \| \, \partial_{x}^2\psi \, \|_{L^2}^{2} \, \mathrm{d}\tau \, \right)^{\frac{1}{2}} \, 
    \left( \, \int^{t}_{0} 
    \| \, \partial_{x}^4\psi \, \|_{L^2}^{2} \, \mathrm{d}\tau \, \right)^{\frac{1}{2}}
\le C_{\phi_0}, 
\end{equation}
for $t \in [\,0, \, T \,]$. 
From the uniform estimates (3.5) and (3.6), 
we immediately get the {\it a priori} estimate for $\partial_{x}\psi$ as follows. 

\medskip

\noindent
{\bf Proposition 3.4.}\quad {\it
For $0\le q \le 5$, there exists a positive constant $C_{\psi_0}$
such that 
\begin{equation*}
\| \, \partial_{x}\psi(t) \, \|_{L^2}^{2} 
  + \int^{t}_{0} 
\| \, \partial_{x}^3\psi(\tau) \, \|_{L^2}^{2} \, \mathrm{d}\tau
\, \mathrm{d}x\mathrm{d}\tau 
\leq C_{\psi_0} 
\quad \big( \, t \in [\, 0, \, T\, ] \, \big).
\end{equation*}
}

From Propositions 3.1, 3.3-3.4, by using the Sobolev inequality, we have 
the following uniform boundedness of $\psi$ and $\partial_{x}\psi$ as follows. 

\medskip

\noindent
{\bf Lemma 3.5.}\quad {\it
There exists a positive constant $C_{\psi_0}$ such that 
$$
\sup_{x \in \mathbb{R}} \, | \, \psi(t,x) \, |\le C_{\psi_0}, 
\quad \sup_{x \in \mathbb{R}} \, | \, \partial_{x}\psi(t,x) \, |\le C_{\psi_0} 
\quad \big( \, t \in [\, 0, \, T\, ] \, \big).
$$
}


\medskip
By combining Propositions 3.1, 3.3-3.4 and Lemma 3.2, 
we can obtain Theorem 2.3. 
Therefore, in order to complete the proof of Theorem 2.3, we finally prove Proposition 3.3. 

\medskip

{\bf Proof of Proposition 3.3.}
Multiplying the equation in (2.2) by 
$\partial_x^4 \psi$, and integrating the resultant formula 
with respect to $x$, we have 
\begin{equation}
\frac{1}{2} \, \frac{\mathrm{d}}{\mathrm{d}t} \, \| \, \partial_x\psi(t) \, \|_{L^2}^2 
+  \nu \, \| \, \partial_x^4\psi(t) \, \|_{L^2}^{2}  
= - \int^{\infty}_{-\infty} 
   \partial_x^4\psi \, 
   \partial_x
   \bigl( \, f(\psi +\tilde{u}) \, \bigr) 
   \, \mathrm{d}x.
\end{equation}
The right-hand side of (3.7) becomes 
\begin{align}
\begin{aligned}
&- \int^{\infty}_{-\infty} 
   \partial_x^4\psi \, 
   \partial_x
   \bigl( \, f(\psi +\tilde{u}) \, \bigr) 
   \, \mathrm{d}x \\
&=\int^{\infty}_{-\infty} 
   f''(\psi + \tilde{u}) \, | \, \partial_x\psi \, |^2 \, \partial_x^3 \psi
   \, \mathrm{d}x+ \int^{\infty}_{-\infty} 
   f'(\psi +\tilde{u}) \, \partial_x^2 \psi \, \partial_x^3 \psi
   \, \mathrm{d}x\\
&=\int^{\infty}_{-\infty} 
   f''(\psi + \tilde{u}) \, | \, \partial_x\psi \, |^2 \, \partial_x^3 \psi
   \, \mathrm{d}x
   -\frac{1}{2} \, \int^{\infty}_{-\infty} 
   f''(\psi +\tilde{u}) \, \partial_x \psi \, | \, \partial_x^2 \psi \,|^2
   \, \mathrm{d}x.
\end{aligned}
\end{align}
From (3.1), 
by making use of the Cauchy-Schwarz, Sobolev and Young inequalities, and 
integration by parts, 
we can estimate the first term on the right-hand side of (3.8) as follows. 
\begin{align}
\begin{aligned}
&
\left| \, 
\int^{\infty}_{-\infty} 
f''(\psi + \tilde{u}) \, | \, \partial_x\psi \, |^2 \, \partial_x^3 \psi
   \, \mathrm{d}x 
   \, \right|\\
&\le C_q \, \int^{\infty}_{-\infty} 
| \, \partial_x\psi \, |^2 \, | \, \partial_x^3 \psi \, |
   \, \mathrm{d}x
+ C_q \, \int^{\infty}_{-\infty} 
     | \, \phi \, |^q \, | \, \partial_x\psi \, |^2 \, | \, \partial_x^3 \psi \, |
   \, \mathrm{d}x, 
\end{aligned}
\end{align}
\begin{align}
\begin{aligned}
&C_q \, \int^{\infty}_{-\infty} 
     | \, \partial_x\psi \, |^2 \, | \, \partial_x^3 \psi \, |
   \, \mathrm{d}x\\
&\le C_q \, \| \, \partial_x\psi \, \|_{L^2}^2 
     \, \| \, \partial_x^3\psi \, \|_{L^2}^{\frac{1}{2}}
     \, \| \, \partial_x^4\psi \, \|_{L^2}^{\frac{1}{2}}\\
&\le C_q \, \| \, \psi \, \|_{L^2} 
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{\frac{5}{4}}
     \, \| \, \partial_x^4\psi \, \|_{L^2}^{\frac{3}{4}}\\
&\le \frac{\nu}{8} \, \| \, \partial_x^4\psi \, \|_{L^2}^2 
     + C_{q, \nu} \, \| \, \psi \, \|_{L^2}^{\frac{8}{5}} 
       \, \| \, \partial_x^2\psi \, \|_{L^2}^2, 
\end{aligned}
\end{align}
\begin{align}
\begin{aligned}
&C_q \, \int^{\infty}_{-\infty} 
     | \, \psi \, |^q \, | \, \partial_x\psi \, |^2 \, | \, \partial_x^3 \psi \, |
   \, \mathrm{d}x\\
&\le C_q \, \| \, \psi \, \|_{L^{\infty}}^{q}
     \, \| \, \partial_x\psi \, \|_{L^2}^2
     \, \| \, \partial_x^3\psi \, \|_{L^{\infty}}\\
&\le C_q \, \| \, \psi \, \|_{L^{\infty}}^{q}
     \, \| \, \psi \, \|_{L^2}
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{\frac{5}{4}}
     \, \| \, \partial_x^4\psi \, \|_{L^2}^{\frac{3}{4}}\\
&\le \frac{\nu}{8} \, \| \, \partial_x^4\psi \, \|_{L^2}^2 
+ C_{q, \nu} \, \| \, \psi \, \|_{L^{\infty}}^{\frac{8q}{5}}
     \, \| \, \psi \, \|_{L^2}^{\frac{8}{5}}
     \, \| \, \partial_x^2\psi \, \|_{L^2}^2\\
&\le \frac{\nu}{8} \, \| \, \partial_x^4\psi \, \|_{L^2}^2 
+ C_{q, \nu} \, \, \left( \, 1+ \| \, \psi \, \|_{L^{\infty}}^8 \, \right)
     \| \, \psi \, \|_{L^2}^{\frac{8}{5}}
     \, \| \, \partial_x^2\psi \, \|_{L^2}^2
     \quad \big( \, 0 \le q \le 5 \, \big).
\end{aligned}
\end{align}
Similarly, the second term on the right-hand side of (3.8) can be estimated as follows. 
\begin{align}
\begin{aligned}
&\left| \, 
\frac{1}{2} \, \int^{\infty}_{-\infty} 
   f''(\psi +\tilde{u}) \, \partial_x \psi \, | \, \partial_x^2 \psi \,|^2
   \, \mathrm{d}x
\, \right|\\
&\le C_q \, \int^{\infty}_{-\infty} 
     | \, \partial_x \psi \, | \, | \, \partial_x^2 \psi \, |^2
   \, \mathrm{d}x
+ C_q \, \int^{\infty}_{-\infty} 
     | \, \psi \, |^{q} \, | \, \partial_x \psi \, | \, | \, \partial_x^2 \psi \, |^2
   \, \mathrm{d}x,
\end{aligned}
\end{align}
\begin{align}
\begin{aligned}
C_q \, \int^{\infty}_{-\infty} 
     | \, \partial_x \psi \, | \, | \, \partial_x^2 \psi \, |^2
   \, \mathrm{d}x
&\le C_q \, \| \, \partial_x\psi \, \|_{L^2}^{\frac{1}{2}} 
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{\frac{5}{2}}\\
&\le C_q \, \| \, \psi \, \|_{L^2}^{\frac{1}{4}} 
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{\frac{11}{4}}\\
&\le C_q \, \| \, \psi \, \|_{L^2}^{\frac{1}{4}} 
     \, \| \, \partial_x^2\psi \, \|_{L^2}^2
     \, \left( \, 1+ \| \, \partial_x^2\psi \, \|_{L^2}^2 \, \right), 
\end{aligned}
\end{align}
\begin{align}
\begin{aligned}
&
C_q \, \int^{\infty}_{-\infty} 
     | \, \psi \, |^{q} \, | \, \partial_x \psi \, | \, | \, \partial_x^2 \psi \, |^2
   \, \mathrm{d}x\\
&\le C_q \, \| \, \psi \, \|_{L^{\infty}}^{q} 
     \, \| \, \partial_x\psi \, \|_{L^{\infty}} 
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{2}\\
&\le C_q \, \left( \, \| \, \psi \, \|_{L^{\infty}}^{\frac{8q}{5}}
     + \| \, \partial_x\psi \, \|_{L^{\infty}}^{\frac{8}{3}} \, \right)
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{2}\\
&\le C_q \, \left( \, 1+ \| \, \psi \, \|_{L^{\infty}}^8
     + \| \, \partial_x\psi \, \|_{L^{\infty}}^{\frac{8}{3}} \, \right)
     \, \| \, \partial_x^2\psi \, \|_{L^2}^{2}
     \quad \big( \, 0 \le q \le 5 \, \big).
\end{aligned}
\end{align}
Noting Lemma 3.2, 
substituting (3.8)-(3.14) into (3.7), 
integrating the resultant formula with respect to $t$ 
and further using the Gronwall inequality, 
we obtain the desired formula, Proposition 3.3. 

Thus, we complete the proof of Theorem 2.3 from Propositions 3.1, 3.3-3.4. 

\bigskip









\bibliographystyle{model6-num-names}
\bibliography{<your-bib-database>}







\end{document}